OpenStudy (countrygirl1431):

Identify this conic section. 16y = x 2 line circle ellipse parabola hyperbola

OpenStudy (igreen):

Here's a graph of it. Can you tell what it is now?

OpenStudy (anonymous):

can you clarify its x^2 or 2.x? \[\ line \ is \ ax + by + x = 0 \\ parabola \ y = ax^2 \\ ellipse\ \frac{x^2}{a^2} + \frac{y^2}{b^2} = 1 \\ hyperbola \frac{x^2}{a^2} - \frac{y^2}{b^2} = 1\]

OpenStudy (anonymous):

\[circle \ (x - a)^2 + (y - b)^2 = r^2\]

OpenStudy (anonymous):

Now you have generalize equation, now think which fit the given equation.

OpenStudy (countrygirl1431):

Parabola?